OpenStudy (anonymous):

the square root of 3t+1 = the square root of t+9

OpenStudy (fortytherapper):

\[\sqrt{3t+1} = \sqrt{t+9}\] Are we doing this problem?

OpenStudy (isaiah.feynman):

Solving for t I'm guessing.

OpenStudy (anonymous):

\[3 t + 1 = t + 9 \] t = 4

OpenStudy (anonymous):

Square each side.

OpenStudy (solomonzelman):

Similar example, \(\color{#000000}{ \displaystyle \sqrt{2x+5}=\sqrt{5x-1} }\) Take the square on both sides (raise both sides to the second power) \(\color{#000000}{ \displaystyle \color{red}{\left(\color{black}{\sqrt{2x+5}}\right)^2} =\color{red}{\left(\color{black}{\sqrt{5x-1}}\right)^2} }\) There is a rule, \(\color{#000000}{ \displaystyle (\sqrt[n]{z})^n=z }\) (and same way, \(\color{#000000}{ \displaystyle (\sqrt[2]{z})^2=z }\)) so, in this case you get, \(\color{#000000}{ \displaystyle 2x+5=5x-1 }\) From now you can solve this arithmetically. \(\color{#000000}{ \displaystyle 2x+5\color{red}{-2x}=5x-1\color{red}{-2x} }\) \(\color{#000000}{ \displaystyle 5=3x-1 }\) \(\color{#000000}{ \displaystyle 5\color{red}{+1}=3x-1\color{red}{+1} }\) \(\color{#000000}{ \displaystyle 6=3x }\) \(\color{#000000}{ \displaystyle 6\color{red}{\div 3}=3x\color{red}{\div 3} }\) \(\color{#000000}{ \displaystyle 2=x }\) If you want/need to check, \(\color{#000000}{ \displaystyle \sqrt{2x+5}=\sqrt{5x-1} }\) \(\color{#000000}{ \displaystyle \sqrt{2\color{red}{(2)} +5}=\sqrt{5\color{red}{(2)}-1} }\) \(\color{#000000}{ \displaystyle \sqrt{4+5}=\sqrt{10-1} }\) \(\color{#000000}{ \displaystyle \sqrt{9}=\sqrt{9} }\)

OpenStudy (anonymous):

\[\sqrt{3t+1} = \sqrt{t+9}\]